The present exchange rate is 4,55 PLN per EUR. You expect the Polish zloty will trade spot at 4,45 in three months. You have 1.000 PLN and you can buy either a put or a call (euro option) at 4,60 PLN for PLN 0,02/EUR. What will you do and what will be your profit if you are correct?
Answer #1
The option price+premium for Euro is more than the
expected spot for Euro after 3 months. Hence, it would be
profitable to buy a put option with the 1000 PLN that is in
hand.
Number of Euros that can be bought under put option = 1000/0.02 = 50000
Amount receivable against the put option = 50000*4.6 = 230000 PLN
Less:
Option premium 1000 PLN
Amount payable in the spot market for 50000 Euros = 50000*4.45 = 222500 PLN
Net profit 6500 PLN